ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 98, 1997

Adams, J. Character of the oscillator representation  229

Bainov, D., Yu. I. Domshlak and P. S. Simeonov On the oscillation properties of first-order impulsive differential equations with a deviating argument

167

Bogdanović, S. See M. Ćirić

Brodsky, S. D. and J. Howie The universal torsion-free image of a group

209

Brozovic, D. P. and R. M. Solomon On groups of hyperbolic length

61

Ćirić, M. and S. Bogdanović The lattice of positive quasi-orders on a semigroup

157

De Sam Lazaro, J. See M. Lemańczyk

deLaubenfels, R., Z. Huang, S. Wang and Y. Wang Laplace transforms of polynomially bounded vector-valued functions and semigroups of operators

189

Domshlak, Yu. I. See D. Bainov

Durdević, M. Generalized braided quantum groups  329

Forrest, A. H. K-groups associated with substitution minimal systems

101

Howie, J. See S. D. Brodsky

Huang, Z. See R. deLaubenfels

Klemes, I. and K. Reinhold Rank one transformations with singular spectral type

1

Kudla, S. S. and W. J. Sweet, Jr. Degenerate principal series representations for U(n, n)

253

Lemańczyk, M. and J. de Sam Lazaro Spectral analysis of certain compact factors for Gaussian dynamical systems

307

Liardet, P. and D. Volný Sums of continuous and differentiable functions in dynamical systems

29

Maharam, D. Spectral representation of generalized transition kernels 15 Mochizuki, K. and R. Suzuki Critical exponent and critical blow-up

for quasilinear parabolic equations

141

Reinhold, K. See I. Klemes Simeonov, P. S. See D. Bainov Solomon, R. M. See D. P. Brozovic Suzuki, R. See K. Mochizuki Sweet, W. J. Jr. See S. S. Kudla Volný, D. See P. Liardet Wang, S. See R. deLaubenfels Wang, Y. See R. deLaubenfels